OpenStudy (anonymous):

u=pa+(p-7) solve for p? sorry these confuses me :/

jimthompson5910 (jim_thompson5910):

u=pa+(p-7) u=pa+p-7 u+7= pa+p u+7= p(a+1) ... ... ... p = ??

OpenStudy (anonymous):

u+7/a+1 ?

jimthompson5910 (jim_thompson5910):

\[\large p = \frac{u+7}{a+1}\] yep